OpenStudy (anonymous):

F(X)=X+3,G(Y)=-6SQUARE ROOT OF Y-1 G(F(X))= -6 SQUARE ROOT OF X+3-1 WHAT IS THE DOMAIN OF THE FUNCTION G(F(X))

OpenStudy (anonymous):

Are these the equations? \[f(x) = x + 3\] \[g(y) = -6\sqrt{y - 1}\]

OpenStudy (anonymous):

YUP YUP

OpenStudy (anonymous):

THE ANSWERS ARE ONLY WORKS FOR EXAMPLE ONE OF THEM IS ALL NUMBERS LESS THAN OR EQUAL TO -3

OpenStudy (anonymous):

No. What I'm asking is if thats what you were trying to write in the question... Is so then \[-6\sqrt{x+2}\]

OpenStudy (anonymous):

And don't you mean -2?

OpenStudy (anonymous):

x + 3 - 1 = x +2

OpenStudy (anonymous):

\[-6\sqrt{X+3-1}\]

OpenStudy (anonymous):

Right! So your domain would equal x ≥ -2

OpenStudy (anonymous):

"x" is greater than or equal to - 2. http://www.wolframalpha.com/input/?i=-6+sqrt%28x+%2B3+-+1%29

OpenStudy (anonymous):

I THINK U DID IT WRONG CUZ THERES NO ANSWERS LIKE THAT THE ONLY NUMVERS WITH THE EQUATIONS ARE -1 AND -3

OpenStudy (anonymous):

Did you look at the link? Thats the equation you provided....

OpenStudy (anonymous):

If anything. Maybe you mean \[x > -3 \] NOT \[x ≥ -3\]

OpenStudy (anonymous):

The lowest domain is -2. \[x ≥ -2\]

OpenStudy (anonymous):

\[F(X)=X+3, G(Y)=-6 \sqrt{Y-1} G(F(X))=\sqrt{X+3-1}\]

OpenStudy (anonymous):

THATS THE PROBLEM

OpenStudy (anonymous):

Oh! I see now... You mean http://www.wolframalpha.com/input/?i=-6+sqrt%28x+%2B3%29+-+1 \[-6\sqrt{x+3} - 1\]

OpenStudy (anonymous):

YEAH

OpenStudy (anonymous):

ONLY THIS WAY, your answers would look like this: Domain: \[x ≥ -3\] Range: \[y ≤ -1\]

OpenStudy (anonymous):

all numbers lees than o equal t0 -3?

OpenStudy (anonymous):

i mean all number less than or equal to -3 would ne my answer

OpenStudy (anonymous):

The originals must have looked like... \[f(x) = x + 3\] \[g(y) = -6\sqrt{y} -1 \]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

The domain goes up. It can only increase. "All real numbers greater than or equal to -3.

OpenStudy (anonymous):

yeah thats what i was thinking of

OpenStudy (anonymous):

so we got it right

OpenStudy (anonymous):

so it would be all numbers greater than or equal to -3

OpenStudy (anonymous):

Yes sir!

OpenStudy (mathstudent55):

\(f(x)=x+3\) \(g(y)=-6 \sqrt{y-1} \) \(g(f(x))= -6 \sqrt{x+3-1} \) \(g(f(x)) = -6\sqrt{x + 2} \) Domain of g(f(x)) = ? The radicand can't be negative, so we solve this equation to find the domain: \(x + 2 \ge 0 \) \(x \ge -2\) That's the domain.

OpenStudy (anonymous):

I'm uncertain about what the equation was... But @mathstudent55 thats what I did above.

OpenStudy (mathstudent55):

Not, exactly. In the composition of functions, you have -1 outside the root.

OpenStudy (anonymous):

Not that one. I mean the one at the very top because he said the options contained only -3 and -1

OpenStudy (mathstudent55):

I see it now. You did get it.

OpenStudy (anonymous):

Thanks! :)